The number of neutrons in the dispositive zinc ion is: (mass number of Zn=65)
A) 35
B) 33
C) 65
D) 67

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Hint: The zinc loses its two electrons and forms dipositive zinc Zn2+ .The atomic number would be equal to the number of protons in the nucleus. The mass number ‘A’ of an atom is equal to the total number of protons and neutrons present in the nucleus of an atom of an element. Complete step by step answer: The atomic number of an atom is the number of protons that are found in the nucleus of the atom of an element. It is denoted by the ‘Z’. Atomic number (Z) = number of proton in nucleus The atomic mass of an atom is defined as the total mass which an atom of any element can exhibit. It is generally expressed in terms of the atomic mass unit. We know that the atoms are made up of the proton, neutrons, and electrons. The nucleus of the atom holds the mass of the atom. The nucleus of an atom is made of the neutrons and protons. The electrons which are negatively charged atomic particles revolve around the nucleus. The atomic mass of the atom is the combined mass of the neutrons and the protons in the nucleus. The mass of electrons is considered negligible concerning the mass of protons and neutrons. Thus, the atomic mass ‘A’ is equal to the sum of neutrons and protons in the atom. Atomic mass (A) = No.of proton + No.of neutron We have given that, The dipositive zinc Zn2+ has the mass number of the 65 (A = 65 ) . We have to find the number of neutrons in the dipositive zinc ion. The zinc is the transition metal. The atomic number of the zinc is equal to the 30. We know that the atomic number of the atom of an element is always equal to the number of protons in the nucleus. Thus equation (1) becomes, Atomic number = Number of proton(p) = 30 Now, put the value of the mass number and the number of proton in the equation (2) we have, Atomic mass (A) = No.of proton + No.of neutron That is, 65 = 30 + No.of neutron∴No.of neutron = 65 −30 = 35 Thus, here we get the total number of neutrons in dipositive zinc ions having a mass of 65 is equal to 35. So, the correct answer is “Option C”. Note: Note that, in zero charge state (neutral state) the number of electrons in the atom of an element is equal to the number of the proton. Atomic number (Z) = number of proton in nucleus = number of electron However, if the atoms are in the charged state (having a positive or negative charge) then the atoms contain less or excess number of electrons in it. In such a case, the above equation is not valid. The Z is equal to the number of protons only. Zn → Zn2+ + 2e− The total number of electrons in the dipositive zinc ion is 28.